OpenStudy (anonymous):

You have 10.0 ml of a 5.0% NaCl solution. To this you add 5.00ml of a 2.0% NaCl solution. What is the molarity of NaCl in this final mixture?

OpenStudy (cwrw238):

you have 0.5 g naCl in 10 ml and 0.1 g in 5 ml so there is 0.6g in 15 mls of solution 1000 mls of a molar solution has 23+ 35.5 = 58.5 g of nacl can you continue from here?

OpenStudy (cwrw238):

you need to calculate the number of grams / litre there are in 1000 mls of the '15 mls' solution = 1000/15 * 0.6 = 40 g so molarity of the solution is 40/58.5 molar

OpenStudy (cwrw238):

= 0.684 molar solution

OpenStudy (anonymous):

Oh gosh. I am so lost in this class :( Thank you so much for everything. I honestly don't even know how you got the .5g of NaCl in 10ml and .1g in 5ml :(

OpenStudy (anonymous):

Oh, never mind...I do know how you got that. Thanks again!

OpenStudy (anonymous):

Where is the "1000 mls" coming from again?